OpenStudy (anonymous):

how to find the laplace of the H(1-t).

OpenStudy (irishboy123):

\[\int\limits_{0}^{ \infty } e ^ {-st} H (t-1) dt = \int\limits_{0}^{1} e ^ {-st} \times 0 \ dt + \int\limits_{1}^{\infty} e ^ {-st} \times 1 \ dt\]

OpenStudy (anonymous):

can u explain how u take 0 and 1.

OpenStudy (irishboy123):

"can u explain how u take 0 and 1." not totally sure i get you exact question but firstly this is the transform of the H function on its own. secondly, H(t-1) is zero until t = 1 and then H = 1 thereafter. so you can just split the interval of integration. hope that helps

OpenStudy (anonymous):

my question is H( 1-t) not H(t-1).

OpenStudy (irishboy123):

crumbs, i see :( H (-x) = 1 - H(x) H(1-t) = 1 - H(t-1) http://math.stackexchange.com/questions/122604/negated-argument-of-the-heaviside-step-function so switch the 0 and the 1 around in my incorrect solution. i now get ( - e^(-s) + 1 ) / s 

OpenStudy (anonymous):

thank u so much.